OpenStudy (anonymous):

Which se contains the zeroes of f(x)= x/x+4-3/x-1/2 A. -6,-4 B. -12,2 C. 6,4 D. 12,-2

OpenStudy (anonymous):

\(f(x)=\frac{x}{x+4} - \frac{3}{x-\frac{1}{2}}\) ?